OpenStudy (anonymous):

Solve for n: 3n+2/7 = 1 - 2n-3/4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol @not-a-cop

OpenStudy (austinl):

First you need to get all the terms with "n" in them on one side.

OpenStudy (anonymous):

hope u can do now ....ahead .....

OpenStudy (anonymous):

yea dog i agree

OpenStudy (anonymous):

This question is also multiple choice, here are the possible answers, A)-1/26 B)41/26 C)10/13 D)7/13

OpenStudy (anonymous):

i think ur options r wrong ... well wait @whpalmer4 @austinL

OpenStudy (anonymous):

yea that's why this make no sense to me @@shkrina

OpenStudy (anonymous):

So, what should I pick for the answer? Even though none of them make sense.

OpenStudy (whpalmer4):

I'm not at all convinced the options are wrong, I think the problem has likely been miswritten here...

OpenStudy (whpalmer4):

@truu can you draw the problem exactly as it appears on your screen, paper, book, stone tablet, whatever?

OpenStudy (austinl):

When I solve out the answer I get -1/140 which is not an answer that is offered. I am of the same belief as @whpalmer4

OpenStudy (anonymous):

sure @whpalmer4

OpenStudy (whpalmer4):

Yeah, this the problem, I bet: \[\frac{3n+2}{7} = 1-\frac{2n-3}{4}\]

OpenStudy (whpalmer4):

The solution to that equation does appear in the list...

OpenStudy (anonymous):

Solve for n: \[\frac{ 3n+2 }{ 7 }=1-\frac{ 2n-3 }{ 4 }\]

OpenStudy (austinl):

That's better!

OpenStudy (anonymous):

girl ur question was wrong .... ,any ways ....

OpenStudy (whpalmer4):

Yeah, okay, in the future, please give us the correct problem at the outset :-) If you can't use the equation editor for the problem statement, use parentheses. \[(3n+2)/7 = 1 - (2n-3)/4\]

OpenStudy (anonymous):

sorry about that @shkrina

OpenStudy (whpalmer4):

In general, just copying down all the symbols won't get you the same meaning, because position matters as well!

OpenStudy (anonymous):

and ok @whpalmer4

OpenStudy (anonymous):

its kk hope u can do it ... by taking LCM and then cross multiply ....

OpenStudy (whpalmer4):

You aren't the first person to make this mistake on OpenStudy, you won't be the last, but maybe it can be the last time you do it :-)

OpenStudy (whpalmer4):

@shkrina even easier would be to multiply by all of the denominators...

OpenStudy (anonymous):

yea I wont mess up other equations again lol @whpalmer4

OpenStudy (anonymous):

1- is their

OpenStudy (whpalmer4):

"is there" :-) that's okay: \[\frac{3n+2}{7} = 1 - \frac{2n-3}{4}\]Now multiply by 7*4 = 28 \[\frac{28*(3n+2)}{7} = 28*1 - \frac{28(2n-3)}{4}\]\[4(3n+2) = 28-7(2n-3)\]\[12n+8 = 28 - 14n + 21\]

OpenStudy (whpalmer4):

poof! no more fractions :-)

OpenStudy (anonymous):

well that is same as LCM ////////

OpenStudy (whpalmer4):

well, with fewer steps :-)

OpenStudy (whpalmer4):

for one, you don't actually have to find the LCM (though numbers may be larger if the denominators are not relatively prime).

OpenStudy (whpalmer4):

@Truu are you able to finish solving that to find the answer?

OpenStudy (anonymous):

I'm kinda confused actually

OpenStudy (whpalmer4):

about how I got there, or how to solve that remaining equation?

OpenStudy (anonymous):

how to solve the remaining, I understand how you got there

OpenStudy (whpalmer4):

\[12n+8=28−14n+21\]Let's get all the \(n\) terms on one side. Add \(14n\) to both sides to get rid of the term on the right side: \[12n + 8 + 14n = 28 -14n + 21 + 14n\]\[26n + 8 = 28 + 21\]Now let's move that 8 to the right side by subtracting 8 from both sides\[26n+8-8=28+21-8\]\[26n = 41\]Now divide both sides by 26 \[\frac{26n}{26} = \frac{41}{26}\]\[n = \frac{41}{26}\]

OpenStudy (whpalmer4):

Hope that's clear enough, I have to go!

OpenStudy (anonymous):

ok yea I knew how to do some of that. Thank you for the help.